# Quantum Computing for Energy Correlators

### v1

@@ Abstract
In recent years, energy correlators have emerged as powerful observables for probing the fragmentation dynamics of high-energy collisions. We introduce the first numerical strategy for calculating energy correlators using the Hamiltonian lattice approach, providing access to the intriguing nonperturbative dynamics of these observables. Furthermore, motivated by rapid advances in quantum computing hardware and algorithms, we propose a quantum algorithm for calculating energy correlators in quantum field theories. This algorithm includes ground state preparation, the application of source, sink, energy flux, real-time evolution operators, and the Hadamard test. We validate our approach by applying it to the SU(2) pure gauge theory in $2+1$ dimensions on $3\times 3$ and $5\times 5$ honeycomb lattices with $j_{\rm max}=\frac{1}{2}$ at various couplings, utilizing both classical methods and the quantum algorithm, the latter tested using the IBM emulator for specific configurations. The results are consistent with the expected behavior of the strong coupling regime and motivate a more comprehensive study to probe the confinement dynamics across the weak and strong coupling regimes.
@@ I Introduction
The energy flow operator in Eq. (1) and its correlations have been widely studied in CFTs [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 11, 17, 18, 19, 20]. Notably, this has led to an improved understanding of the operator product expansion (OPE) for Lorentzian operators, specifically the light-ray OPE [21, 22, 23]. In recent years, there has been a strong motivation to measure such light-ray OPE in collider experiments by studying the substructure of jets [24, 25], where the energy flow operator in Eq. (1) represents the experimental detectors that capture the energy flux. (For this reason, we will use ‘energy flow operator’ and ‘detector’ interchangeably.) This has resulted in the observation of universal scaling predicted by the light-ray OPE across multiple experimental collaborations [26, 27, 28], along with a surge in phenomenological studies using energy correlators to explore the Lorentzian dynamics of quantum chromodynamics (QCD) [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61].
@@
In recent years, the Hamiltonian lattice approach emerged as a promising tool for nonperturbative studies of the Lorentzian dynamics of quantum field theories (QFTs), driven by advances in quantum computing hardware and algorithms, as well as classical techniques for real-time dynamics [66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97]. Recent progress includes the development of classical and quantum algorithms for simulating gauge theories, which are central to our understanding of high-energy physics. Notably, there have been successful demonstrations of classical and quantum simulations of lattice gauge theories in lower dimensions, providing a foundation for exploring more intricate theories relevant to QCD [98, 99, 100, 101, 102, 67, 103, 104, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115].
In this paper, we capitalize on these advancements and develop a quantum algorithm to compute energy correlators nonperturbatively for a generic QFT using the Hadamard test [116]. As a demonstration, we consider the $2+1$-dimensional SU(2) pure gauge theory, a confining theory [117, 118, 119, 120, 121, 122], and calculate energy correlators on a small lattice with fixed couplings and truncated local Hilbert spaces. We apply the quantum algorithm to a $3\times 3$ honeycomb lattice and classically simulate it on both $3\times 3$ and $5\times 5$ honeycomb lattices as schematically shown in Fig. 1. Our study lays the groundwork for future calculations in the continuum limit and demonstrates how quantum computing can be valuable for high-energy physics, even without the complicated initial wave packet preparation typically required for scattering simulations [123, 124], although progress has been made to prepare wave packets [102, 101].
@@ II.1.1 State created by local source
We first consider the case where both $\mathcal{O}$ and $\mathcal{O}^{\dagger}$ are local operators inserted at the same spacetime point $x$, i.e., $\mathcal{O}=\mathcal{O}(x)=\mathcal{O}(t,\vec{x})$, $\mathcal{O}^{\dagger}=\mathcal{O}^{\dagger}(x)=\mathcal{O}^{\dagger}(t,\vec{x})$. The last terms spell out the time and spatial components. Without loss of generality, we assume they are placed at the spacetime origin $x=0$, denoted by the red point in Fig. 2. To compute the energy correlators, we employ a detector limiting procedure as in Eq. (1), i.e. light transform [23, 12, 126, 127], by sending the stress-energy tensor $T_{0i}$ to infinity and performing the time integration. This procedure represents constructing an asymptotic detector which turn on during some detector operation time.
Figure 2: Penrose diagram for $d+1$ dimensional Minkowski space depicting the integration and limit procedure for constructing an energy flow operator. Each point represents $d-1$ dimensional sphere. The Penrose coordinates $\tilde{t}$ and $\tilde{r}$ are related to the Minkowski coordinates $t$ and $r$ as $\tan(\tilde{t}\pm\tilde{r})=t\pm r=x^{\pm}$, and manifest the causal structure of the spacetime. The red point represents a local source of perturbation from which the signals travel. The two dashed red lines indicate different paths taken by massive and massless excitations from the source at $(t,r)=(0,0)$, reaching timelike infinity $i^{+}$ and null future infinity $\mathscr{I}^{+}$, respectively. The $i^{0}$ indicates spacelike infinity. The solid blue line represents the continuum integration path of the detector that is turned on at an initial position $r_{d}$ from $t=0$ to $t_{d}$ with $x^{+}$ constant. The dashed curved blue line approximates this path for lattice calculations. We take the $x^{+}\to\infty$ limit by taking $r_{d}$ to $i^{0}$ and $t_{d}$ to $i^{+}$ while moving the blue line along the dotted blue arrow line.
@@
where the integrated correlation $G_{{\vec{n}}_{1}{\vec{n}}_{2}}(x^{+})$ is at a fixed advanced time. We have used the fact that the QFT is in $d+1$ dimensions. We also note that the bound for $x^{-}$ is determined assuming that the source is inserted at $t=0$. This procedure is natural for massless excitations, such as those in CFTs ending in the future null infinity. For gapped theories like the SU(2) gauge theory we consider later, the integration over retarded time should be taken before sending the advanced time to infinity, in order to capture massive excitations at timelike infinity, as noted in Ref. [128].
Keeping all the detectors at fixed $x^{+}$ while the integration is performed maintains spacelike separation between the detectors on the same light-sheet even for finite $r$. For $\vec{n}_{1}\neq\vec{n}_{2}$, which we focus on this paper, we then expect the energy flow operators to commute [12, 11, 10, 17]
@@
Due to the discretization inherent in the lattice, we must also discretize our detector limiting procedure. We approximate the smooth integration path by a discretized version, as shown by the blue dashed line in the Penrose diagram in Fig. 2. Following this numerical procedure, we can approximate $G_{{\vec{n}}_{1}{\vec{n}}_{2}}(x^{+})$ in Eq. (II.1.1) on the lattice as
@@
We note that at each step $i$, we fix $r_{i}+t_{i}=x^{+}$ as desired. We also take $t_{0}=0$ as we identify $t=0$ with the time we insert the source. In the Hamiltonian lattice theory setup, time is continuous but space is discretized. The smallest step we can then take is given by the lattice spacing $a$, i.e., $r_{0}-r_{N}=Na$. (To reduce the lattice cutoff artifact, one should take each step to be a few lattice spacings, say $Ma$, which would shift $r_{0}-r_{N}=NMa$.) At finite $a$, the relation to keep $x^{+}$ fixed is slightly broken when we integrate between $t_{i}$ and ${t}_{i+1}$ with fixed $r_{i}$. This can spoil the spacelike separation that we want to ensure in Eq. (4), leading to causal interference between detectors. To mitigate this undesired lattice artifact, we set the upper limit of the time integration at each step $i$ to $\tilde{t}_{i}\leq t_{i+1}$, which means we miss energy flux between $\tilde{t}_{i}$ and $t_{i+1}$. However, as the lattice spacing $a\to 0$, these artifacts should vanish, allowing us to take $\tilde{t}_{i}\to t_{i+1}$ in the continuum limit. We will explain our choice of $\tilde{t}_{i}$ later in our explicit setup for the $2+1$ SU(2) theory on the honeycomb lattice.
@@ II.1.2 Momentum eigenstate
In the previous section, we discussed local sources and assumed both $\mathcal{O}$ and $\mathcal{O}^{\dagger}$ to be inserted at the spacetime origin. In this subsection, we discuss a strategy to studying energy correlators of momentum eigenstates which are not localized. For example, in $e^{+}e^{-}$ collisions, the electromagnetic current $J^{\mu}(x)=e\bar{\psi}(x)\gamma^{\mu}\psi(x)$ can source a state with definite energy and momentum $q$, then we can define an energy-energy correlator as (see e.g. Ref. [126])
@@
which is same as Eq. (II.1.2) in the limit. However, for a finite value of $\sigma$, the Gaussian suppression factor ensures that Eq. (II.1.2) has finite support with a width of $\sigma$ around the point $(x_{0},\vec{x})=(0,\vec{0})$. Neglecting small contributions from the Gaussian tail, we can define a time extent $t_{s}$ and a spatial extent $r_{s}$, within which the source and sink terms are confined, as illustrated in Fig. 3. Unlike the scenario in Fig. 2, where the source and sink terms are inserted at a single point, the finite region necessitates placing the detector sufficiently far away to ensure that no signals are missed. Signals traveling the farthest are emitted radially outward from $(-t_{s},r_{s})$ and reach $r=t_{s}+r_{s}$ at $t=0$. Consequently, the detector must be positioned at $r_{d}\geq t_{s}+r_{s}$ as in Fig. 3 to capture all signals from the source. In other words, the width $\sigma$ effectively defines a finite region where the source and sink terms are inserted, which then imposes a lower bound on the distance at which the detector must be placed. We can then follow the implementation in Eq. (II.1.1) with $r_{0}=x^{+}\geq t_{s}+r_{s}$ and an additional integral over the finite region controlled by $\sigma$, where the source and sink terms are inserted, to study energy correlators of the momentum eigenstate in the limit of Eq. (II.1.2).
@@ II.2 Real-time correlator from Hamiltonian dynamics and classical simulation
If we can obtain all the eigenvalues and eigenstates of the system, which are defined as $H|m\rangle=E_{m}|m\rangle$, we can calculate the real time correlators in Eq. (II.2) as
@@ II.3 A quantum algorithm
Eq. (II.2) informs us on how to construct the quantum algorithm to evaluate the real time correlator, which is then integrated to obtain the energy-energy correlator as in Eqs. (II.1.1) and (II.1.1). We will now introduce a quantum algorithm that is based on the Hadamard test [116]. A schematic diagram of the algorithm is shown in Fig. 4, which consists of the following two steps:
@@ II.3.1 Ground state preparation
The first step of the quantum algorithm prepares the ground state $|{\rm vac}\rangle$ of the interacting system by starting from the qubit default state $|000\dots 0\rangle$ [^4] In order to avoid confusion between the interacting vacuum state of the theory $|0\rangle$ and the qubit state $|000\dots 0\rangle$, we have used $|{\rm vac}\rangle$ to label the former in the description of quantum circuits. via e.g., the adiabatic time evolution [129, 130]. The algorithm simulates time evolution driven a time-dependent Hamiltonian
@@
where $H_{0}$ is a Hamiltonian whose ground state is easy to prepare, such as a product state in the computational basis. $H_{0}+H_{1}$ is the target Hamiltonian for which we want to prepare the ground state. The initial state of the time evolution is given by the ground state of $H_{0}$. The run time $T$ of the adiabatic ground state preparation algorithm scales as $T=O(1/[\min\Delta E(t)]^{2})$ where $\Delta E(t)=E_{1}(t)-E_{0}(t)$ is the energy gap between the ground state and the first excited state of the time-dependent Hamiltonian $H(t)$ [131]. For SU(2) and SU(3) non-Abelian gauge theories in two or three spatial dimensions that are confining, we expect a mass gap in the spectrum even in the continuum limit.
@@
One may also use variational methods, such as Variational Quantum Eigensolver (VQE) or ADAPT-VQE [132, 133, 134, 135], to construct a shallow quantum circuit that prepares the ground state or the state |$\ket{\Psi}=\mathcal{O}|{\rm vac}\rangle$ as well. These algorithms are more empirical than the adiabatic approach in the sense that the variational operator ansatz and their parameters depend on the system of interest.

### v2

@@ Abstract
In recent years, energy correlators have emerged as powerful observables for probing the fragmentation dynamics of high-energy collisions. We introduce the first numerical strategy for calculating energy correlators using the Hamiltonian lattice approach, providing access to the intriguing nonperturbative dynamics of these observables. Furthermore, motivated by rapid advances in quantum computing hardware and algorithms, we propose a quantum algorithm for calculating energy correlators in quantum field theories. This algorithm includes ground state preparation, the application of source, sink, energy flux and real-time evolution operators, and the Hadamard test. We validate our approach by applying it to the SU(2) pure gauge theory in $2+1$ dimensions on $3\times 3$ and $5\times 5$ honeycomb lattices with $j_{\rm max}=\frac{1}{2}$ at various couplings, utilizing both classical methods and the quantum algorithm, the latter tested using the IBM emulator for specific configurations. The results are consistent with the expected behavior of the strong coupling regime and motivate a more comprehensive study to probe the confinement dynamics across the weak and strong coupling regimes.
@@ I Introduction
The energy flow operator in Eq. (1) and its correlations have been widely studied in CFTs [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 11, 17, 18, 19, 20]. Notably, this has led to an improved understanding of the operator product expansion (OPE) for Lorentzian operators, specifically the light-ray OPE [21, 22, 23]. In recent years, there has been a strong motivation to measure such light-ray OPE in collider experiments by studying the substructure of jets [24, 25], where the energy flow operator in Eq. (1) represents the experimental detectors that capture the energy flux. (For this reason, we will use ‘energy flow operator’ and ‘detector’ interchangeably.) This has resulted in the observation of universal scaling predicted by the light-ray OPE across multiple experimental collaborations [26, 27, 28], along with a surge in phenomenological studies using energy correlators to explore the Lorentzian dynamics of quantum chromodynamics (QCD) [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62].
@@
In recent years, the Hamiltonian lattice approach emerged as a promising tool for nonperturbative studies of the Lorentzian dynamics of quantum field theories (QFTs), driven by advances in quantum computing hardware and algorithms, as well as classical techniques for real-time dynamics [67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97]. Recent progress includes the development of classical and quantum algorithms for simulating gauge theories, which are central to our understanding of high-energy physics. Notably, there have been successful demonstrations of classical and quantum simulations of lattice gauge theories in lower dimensions, providing a foundation for exploring more intricate theories relevant to QCD [98, 99, 100, 101, 102, 103, 104, 105, 106, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117].
In this paper, we capitalize on these advancements and develop a quantum algorithm to compute energy correlators nonperturbatively for a generic QFT using the Hadamard test [118]. As a demonstration, we consider the $2+1$-dimensional SU(2) pure gauge theory, a confining theory [119, 120, 121, 122, 123, 124], and calculate energy correlators on a small lattice with fixed couplings and truncated local Hilbert spaces. We test the quantum algorithm on a $3\times 3$ honeycomb lattice and classically calculate energy correlators on both $3\times 3$ and $5\times 5$ honeycomb lattices as schematically shown in Fig. 1. Our study lays the groundwork for future calculations in the continuum limit and demonstrates how quantum computing can be valuable for high-energy physics, even without the complicated initial wave packet preparation typically required for scattering simulations [125, 126], although progress has been made to prepare wave packets [103, 102].
@@ II.1.1 State created by local source
We first consider the case where both $\mathcal{O}$ and $\mathcal{O}^{\dagger}$ are local operators inserted at the same spacetime point $x$, i.e., $\mathcal{O}=\mathcal{O}(x)=\mathcal{O}(t,\vec{x})$, $\mathcal{O}^{\dagger}=\mathcal{O}^{\dagger}(x)=\mathcal{O}^{\dagger}(t,\vec{x})$. Without loss of generality, we assume they are placed at the spacetime origin $x=0$, denoted by the red point in Fig. 2. To compute the energy correlators, one needs to employ a detector limiting procedure as in Eq. (1), i.e. light transform [23, 12, 128, 129], by sending the stress-energy tensor $T_{0i}$ to infinity and performing the time integration. This procedure represents constructing an asymptotic detector which turn on during some detector operation time.
Figure 2: Penrose diagram for $d+1$ dimensional Minkowski space depicting the integration and limiting procedure for constructing an energy flow operator. Each point represents $d-1$ dimensional sphere. The Penrose coordinates $\tilde{t}$ and $\tilde{r}$ are related to the Minkowski coordinates $t$ and $r$ as $\tan(\tilde{t}\pm\tilde{r})=t\pm r=x^{\pm}$, and manifest the causal structure of the spacetime. The red point represents a local source of perturbation from which the signals travel. The two red dashed lines indicate different paths taken by massive and massless excitations from the source at $(t,r)=(0,0)$, reaching timelike infinity $i^{+}$ and null future infinity $\mathscr{I}^{+}$, respectively. The $i^{0}$ indicates spacelike infinity. The blue solid line represents the continuum integration path of the detector that is turned on at an initial position $r_{d}$ from $t=0$ to $t_{d}$ with $x^{+}$ constant. The blue dashed curved line approximates this path for lattice calculations. We take the $x^{+}\to\infty$ limit by taking $r_{d}$ to $i^{0}$ and $t_{d}$ to $i^{+}$ while moving the blue solid line along the blue dotted arrow line.
@@
where the integrated correlation $G_{{\vec{n}}_{1}{\vec{n}}_{2}}(x^{+})$ is at a fixed advanced time. We have used the fact that the QFT is in $d+1$ dimensions. We also note that the bound for $x^{-}$ is determined assuming that the source is inserted at $t=0$. This procedure is natural for massless excitations, such as those in CFTs ending in the future null infinity. For gapped theories like the SU(2) gauge theory we consider later, the integration over the retarded time should be taken before sending the advanced time to infinity, in order to capture massive excitations at timelike infinity, as noted in Ref. [130].
Keeping all the detectors at fixed $x^{+}$ while the integration is performed maintains spacelike separation between the detectors on the same light-sheet even for finite $r$. For $\vec{n}_{1}\neq\vec{n}_{2}$, which we focus on in this paper, we then expect the energy flow operators to commute [12, 11, 10, 17]
@@
Due to the discretization inherent in the lattice, we must also discretize our detector limiting procedure. We approximate the smooth integration path with a discretized version, as shown by the blue dashed line in the Penrose diagram in Fig. 2. Following this numerical procedure, we can approximate $G_{{\vec{n}}_{1}{\vec{n}}_{2}}(x^{+})$ in Eq. (II.1.1) on the lattice as
@@
We note that at each step $i$, we fix $r_{i}+t_{i}=x^{+}$ as desired. We also take $t_{0}=0$ as we identify $t=0$ with the time we insert the source. In the Hamiltonian lattice theory setup, time is continuous but space is discretized. The smallest step we can take is given by the lattice spacing $a$, i.e., $r_{0}-r_{N}=Na$. (To reduce the lattice cutoff artifact, one should take each step to be a few lattice spacings, say $Ma$, which would give $r_{0}-r_{N}=NMa$.) At finite $a$, the relation to keep $x^{+}$ fixed is slightly broken when we integrate between $t_{i}$ and ${t}_{i+1}$ with fixed $r_{i}$. This can spoil the spacelike separation that we want to ensure in Eq. (4), leading to causal interference between detectors. To mitigate this undesired lattice artifact, we set the upper limit of the time integration at each step $i$ to $\tilde{t}_{i}\leq t_{i+1}$, which means we miss energy flux between $\tilde{t}_{i}$ and $t_{i+1}$. However, as the lattice spacing $a\to 0$, these artifacts should disappear, allowing us to take $\tilde{t}_{i}\to t_{i+1}$ in the continuum limit. We will explain our choice of $\tilde{t}_{i}$ later in our explicit setup for the $2+1$ SU(2) theory on the honeycomb lattice.
@@ II.1.2 Momentum eigenstate
In the previous section, we discussed local sources and assumed that both $\mathcal{O}$ and $\mathcal{O}^{\dagger}$ were inserted at the origin of spacetime. In this subsection, we introduce a strategy for calculating energy correlators of momentum eigenstates that are not localized. For example, in $e^{+}e^{-}$ collisions, the electromagnetic current $J^{\mu}(x)=e\bar{\psi}(x)\gamma^{\mu}\psi(x)$ can source a state with definite energy and momentum $q$, then we can define an energy-energy correlator as (see e.g. Ref. [128])
@@
which is same as Eq. (II.1.2) in the limit. However, for a finite value of $\sigma$, the Gaussian suppression factor ensures that Eq. (II.1.2) has finite support with a width of $\sigma$ around the point $(t,\vec{x})=(0,\vec{0})$. Neglecting small contributions from the Gaussian tail, we can define a time extent $t_{s}$ and a spatial extent $r_{s}$, within which the source and sink terms are applied, as illustrated in Fig. 3. Unlike the scenario in Fig. 2, where the source and sink terms are inserted at a single point, the finite region necessitates placing the detector sufficiently far away to ensure that no signals are missed. Signals traveling the farthest are emitted radially outward from $(-t_{s},r_{s})$ and reach $r=t_{s}+r_{s}$ at $t=0$. Consequently, the detector must be positioned at $r_{d}\geq t_{s}+r_{s}$ as in Fig. 3 to capture all signals from the source. In other words, the width $\sigma$ effectively defines a finite region where the source and sink terms are inserted, which then imposes a lower bound on the distance at which the detector must be placed. We can then follow the implementation in Eq. (II.1.1) with $r_{0}=x^{+}\geq t_{s}+r_{s}$ and an additional integral over the finite region controlled by $\sigma$, where the source and sink terms are inserted, to study energy correlators of the momentum eigenstate in the limit of Eq. (II.1.2).
@@ II.2 Real-time correlator from Hamiltonian dynamics and classical simulation
If we can obtain all the eigenvalues and eigenstates of the system, which are defined as $H|m\rangle=E_{m}|m\rangle$, we can calculate the real-time correlators in Eq. (II.2) as
@@ II.3 A quantum algorithm
Eq. (II.2) informs us on how to construct the quantum algorithm to evaluate the real-time correlator, which is then integrated to obtain the energy-energy correlator as in Eqs. (II.1.1) and (II.1.1). We will now introduce a quantum algorithm that is based on the Hadamard test [118]. A schematic diagram of the algorithm is shown in Fig. 4, which consists of the following two steps:
@@ II.3.1 Ground state preparation
The first step of the quantum algorithm prepares the ground state $|{\rm vac}\rangle$ of the interacting system by starting from the qubit default state $|000\dots 0\rangle$ [^4] In order to avoid confusion between the interacting vacuum state of the theory $|0\rangle$ and the qubit state $|000\dots 0\rangle$, we use $|{\rm vac}\rangle$ to label the former in the description of quantum circuits. via e.g., the adiabatic time evolution [131, 132]. The algorithm simulates the time evolution driven by a time-dependent Hamiltonian
@@
where $H_{0}$ is a Hamiltonian whose ground state is easy to prepare, such as a product state in the computational basis. $H_{0}+H_{1}$ is the target Hamiltonian for which we want to prepare the ground state. The initial state of the time evolution is given by the ground state of $H_{0}$. The run time $T$ of the adiabatic ground state preparation algorithm scales as $T=O(1/[\min\Delta E(t)]^{2})$ where $\Delta E(t)=E_{1}(t)-E_{0}(t)$ is the energy gap between the ground state and the first excited state of the time-dependent Hamiltonian $H_{\rm ad}(t)$ [133]. For SU(2) and SU(3) non-Abelian gauge theories in two or three spatial dimensions that are confining, we expect a mass gap in the spectrum even in the continuum limit.
@@
One may also use variational methods, such as the Variational Quantum Eigensolver (VQE) or ADAPT-VQE [134, 135, 136, 137], to construct a shallow quantum circuit that prepares the ground state or the state |$\ket{\Psi}=\mathcal{O}|{\rm vac}\rangle$ as well. These algorithms are more empirical than the adiabatic approach in the sense that the variational operator ansatz and their parameters depend on the system details and may need reconstructing at different system sizes and couplings, particularly for the continuum extrapolation.